How can this concept be extended to systems involving three variables and three equations as well as solving the following system of three equations and describing the methods you are using:

X + Y + Z = 6
2X – Y + 3Z = 8
3X – 2Y – Z = -17

Answer by Alan3354(69443)

How can this concept be extended to systems involving three variables and three equations as well as solving the following system of three equations and describing the methods you are using:

X + Y + Z = 6
2X – Y + 3Z = 8
3X – 2Y – Z = -17
===================
What concept?